OpenStudy (anonymous):

All of the following are equivalent except ___? 33 1/3% 0.3^- 1/3 33.3^-

OpenStudy (anonymous):

what is 0.3^- and 33.3^- ?

OpenStudy (anonymous):

33.3333.....

OpenStudy (anonymous):

let x=0.333.... ...(1) 10x=3.333... ...(2) (2)-(1) gives 9x=3.00 x=3/9=1/3

OpenStudy (anonymous):

33 1/3 %=100/3*1/100=1/3 options 1,2,3 are equivalent